OpenStudy (anonymous):

probability question ? what is probability of getting head first time head second time and head third time? in fair coin.

Parth (parthkohli):

These are three INDEPENDENT EVENTS, and to find the probability of these happening together, you must multiply the probability of each.

Parth (parthkohli):

P(head first time) = 1/2 P(head second time) = 1/2 P(head third time) = 1/2 P(all three) = 1/2 * 1/2 * 1/2

OpenStudy (campbell_st):

its simply P(head) x P(head) x P(head) or draw a free diagram for the experiment

Parth (parthkohli):

Note that what you get on any turn does not depend on what you got on the previous term.

OpenStudy (anonymous):

thanks . read the next equation am posting plz